# Human-In-The-Loop Machine Learning for Safe and Ethical Autonomous Vehicles: Principles, Challenges, and Opportunities

### v1

@@ III-C2 Integration of Human Feedback in HITL-RL
III-C3 Agent-Specific vs. Agent-Agnostic Protocols
Agent-specific protocols, meaning protocols for human interaction that are designed for a specific HITL-RL algorithm are preferred over agent-agnostic protocols. Agent-specific protocols allow humans to provide feedback and advice that is directly relevant to the AV’s learning algorithm. By understanding the vehicle’s specific learning process and representation, the protocols enable humans to ask more precise and informative questions. This leads to higher-quality guidance that is more likely to accelerate the vehicle’s learning and improve its decision-making capabilities. In contrast, agent-agnostic protocols are not specialized to the agent and therefore cannot ask the human informative questions or have an observation model that accurately represents the human’s process for generating advice. Similarly, humans cannot provide optimally informative advice to the agent because they are unaware of the agent’s prior knowledge, exploration techniques, representation, or learning method[70].
III-C4 Evaluating HITL-ML Algorithms
Despite of significant contributions towards HITL-ML algorithms, most studies focus on improving the performance of algorithms rather than the interaction experience, but proper evaluation is essential to create practical applications for different types of end-users. The HITL-ML necessitates incorporating subjective user evaluations and acknowledging the cyclic influence between algorithms and users, making HITL-ML performance evaluation inherently subjective and complex[71, 72].
III-C5 Robustness Against Adversarial Attacks
HITL-ML can make the models more robust against adversarial attacks. Adversarial attacks are deliberate attempts to fool ML models by providing deceptive input. When human are engaged in the model development process, they can provide valuable insights, detect anomalies, and make adjustments that might not be apparent through automated processes alone. This human involvement can help identify and mitigate potential vulnerabilities that adversarial attacks might exploit, thus reducing the severity and impact of such attacks on the models[73],[74].
III-C6 Combining with safe RL Techniques
The addition of Safe RL techniques to HITL-RL setting can further reduce the likely of dangerous actions. Safe RL techniques ensure good performance and respect safety constraints during learning and deployment[73],[75].
Figure 5: The diagram of AL for AVs is highlighted, where AVs face unknown and unexpected scenarios and utilize AL to annotate them. At the corner, the AL loop is depicted and comprises the steps of training, object selection, annotating, and appending. In the annotation step, human experts annotate the selected examples. AL impacts on unseen situtations and training also highlighted.
@@ IV Active Learning
This section introduces the topic and reviews various studies on AL in different domains of AVs, from UAV anomaly mitigation and aerial semantic mapping to intrusion detection and 3D object detection in AD. These applications demonstrate how AL optimizes the annotation process, accelerates AV development, and improves the robustness of vehicle-related systems by effectively leveraging human expertise.
IV-A Preliminaries
We introduce AL in AVs, reviewing studies across various domains. AL optimizes the annotation process, accelerates development, and enhances system robustness by leveraging human expertise effectively. Annotating extensive data for training AVs is challenging due to their complex environments and unexpected situations. AL addresses these issues by reducing manual labeling through targeted instance selection, improving the quality of labeled data, and the accuracy of networks. The amount of data needed to train autonomous aerial or ground vehicles is enormous. AVs must navigate a variety of complex and diverse environments, which necessitate being represented in the training data. Moreover, AVs must deal with unexpected situations. Thus, to master these events, the model needs extensive data that captures such occurrences. Annotating such a volume of data for training autonomous aerial and ground vehicles can be an overwhelming and tedious task for humans and can pose a major challenge to AV development.
AL can be applied to this process, to cut down on the time and cost spent on training, while also increasing the accuracy of the networks. AL makes it possible to automate the selection process while choosing valuable data points. It starts by training a dedicated ML model on already-labeled data. The network then sorts through unlabeled data, uses the acquisition function to identify the instances it’s most uncertain about, thereby finding data that would be challenging to the AVs algorithm and asks a human to label these specific instances which are added to training data and then the model is trained with labeled data. This approach reduces the number of instances that require manual labeling and ensures that the labeled data is as informative as possible [76]. As can be seen in Fig. 5, AVs are confronted with unknown and unexpected scenarios and use AL to annotate them. In the corner, the AL loop is shown, which includes the training, object selection, annotation and appending steps. In the annotation step, the selected examples are annotated by human experts. AL can help to cope with unseen situations, reduce training time and costs and improve training efficiency.
@@ IV-A Preliminaries
IV-B Literature Review
IV-B1 Security and Navigation
The authors of [41] aim to mitigate in-flight UAV anomalies by using a semi-supervised SVM. Since the pure semi-supervised SVM classification method may not classify well due to a lack of unlabeled data with rich content. AL is used to search the samples with rich information and significant differences for labeling to improve the quality of labeled data. Building on this concept, the authors of [42] have developed a new path planning framework for AL in aerial semantic mapping to improve UAV semantic awareness with minimal expert guidance. In an innovative approach, path planning goals are combined with AL features that allow us to adapt the UAV’s path to regions with informative training data. Similarly, the authors of [43] design the intrusion detection system to detect unknown network attacks in the UAV network. The Open-CNN model is proposed to detect unknown attacks. Meanwhile, an AL approach is proposed for the unknown attacks based on the least confidence strategy, which enables the intrusion detection model to efficiently learn about observed novel network attacks with a low labeling budget. Table VIII summarizes the above findings.
IV-B2 Detection
The authors of [47] are concerned with 3D object detection in AV. DL-based object detectors require large annotated datasets, which incurs compilation costs. To address this problem, they developed diversity-based AL, which uses the multimodal information of an AV dataset to propose a novel detection function that enforces spatial and temporal diversity in the selected samples. Simply, the most informative frames and objects are automatically selected for annotation by humans. Similarly, the authors of [15] address vehicle type recognition (VTR) from surveillance data in intelligent transportation. To facilitate manual labeling of large-scale surveillance data, they propose a VTR system based on deep AL. A deep neural network is efficiently trained by actively, incrementally, and automatically selecting unlabeled data points for human labeling. The training procedure is repeated several times so that the neural network is re-trained with the incrementally updated training data. Unlike traditional AL where the samples with high entropy are selected as the network output, in the proposed method, samples with both high and low entropy values are selected to be included in the training data. The proposed method is evaluated using the surveillance images from the CompCars public dataset. In addition, the authors of [77] develop a two-part technique for vehicle detection using an AL system and a symmetry-based iterative analysis. A modified AL technique is proposed that automatically selects positive and negative samples from multiple parts to train adaboost classifiers. The results of the classifiers are linked using a novel iterative window search algorithm and a symmetry-based regression model to extract fully visible vehicles. The AL method combined with the symmetry-based analysis has been shown to improve detection rates and also reduce false alarms compared to existing methods. Table IX summarizes the above findings.
@@ IV-B2 Detection
We have seen the HITL-RL, AL, and CL collectively improve the capabilities of AVs, they also bring forth critical ethical and legal considerations. AVs should act as ethically rational agents and consider the ethical and legal implications of their actions, not just the outcomes.
IV-C Key Findings and Insights
AL streamlines the annotation process by targeting specific instances that need to be labeled, reducing the overall time and cost associated with training AVs. By automating the selection of instances for annotation, AL ensures that the labeled data is maximally informative, improving the efficiency of the training process. AL helps improve the robustness of vehicle-related systems by ensuring that the training data captures diverse and unexpected scenarios. By focusing on labeling instances that are most uncertain for the model, AL helps to overcome the challenge of unexpected situations, ultimately leading to more robust AV systems. Various areas within AV development, such as UAV anomaly mitigation, aerial semantic mapping, intrusion detection, object detection, and vehicle type detection, require tailored AL approaches. Each domain may have unique challenges and requirements, and AL techniques must be adapted accordingly to effectively address these challenges and optimize the training process. AL can be seamlessly integrated into existing techniques and frameworks within AV development to enhance performance and efficiency. By combining AL with other methods, such as DL, SVMs, or symmetry-based analysis, researchers can achieve superior results in tasks such as anomaly detection, semantic mapping, object detection, and vehicle recognition.
Using a semi-supervised SVM with AL significantly improves the quality of labeled data, effectively mitigating in-flight UAV anomalies by selecting samples with rich information and significant differences for labeling. Combining path planning goals with AL features allows UAVs to adapt their paths to regions with informative training data, enhancing semantic awareness with minimal expert guidance. An Open-CNN model combined with an AL approach based on the least confidence strategy efficiently learns about novel network attacks with a low labeling budget, improving the detection of unknown attacks in UAV networks. A deep AL-based VTR system effectively facilitates manual labeling by incrementally selecting high and low-entropy samples for human annotation, significantly improving training efficiency and accuracy using the CompCars dataset. Combining an AL technique with symmetry-based iterative analysis improves vehicle detection rates and reduces false alarms by automatically selecting positive and negative samples and linking classifier results through an iterative window search algorithm and a symmetry-based regression model.
IV-C1 Btach Training VS Sequential Training
@@ IV-C4 Uncertainty Sampling
IV-C5 Diverse Road Condition
@@ V Ethical Implications of AVs
This section reviews the state-of-the-art in ethical AVs and provides a discussion on ethical AVs using HITL-ML, we emphasize that HITL-ML can promote self-recognition and scene recognition in AVs to assist in managing complex situations such as accidents and conducting ethical decisions in real-time.
@@
V-A Preiliminieries
It is of utmost importance to ensure that ML algorithms in AI behave ethically and to design ML algorithms to make ethical decisions. Ethical principles must be encoded into AVs to ensure that AV behavior aligns with societal values and norms. Developing ML algorithms in AVs that can make decisions based on ethical considerations is imperative. To this end, ethical theories and principles must be integrated into the decision-making processes of AVs. AVs should act as ethically rational agents and consider the ethical implications of their actions, not just the outcomes. For example, an autonomous UAV can be considered a rational ethical agent if it accurately perceives its environment and makes decisions based on ethical frameworks such as utilitarianism where the ethical worth of an action is evaluated based on its consequences, with a focus on maximizing utility or happiness[87], or deontology where the ethical actions should be guided by a set of ethical principles or rules, regardless of the consequences [88], and acts in a way that is consistent with ethical principles, ensuring that its actions are “right” and beneficial to all stakeholders.
V-B Literature Review
Complex situations involving human live need ethical consideration. The authors of [78] analyze the legal and ethical aspects of using UAVs, and state that we can achieve a higher degree of trust by formalizing the process of ensuring the safe and compatible functioning of systems based on AI and ML. However, there are some challenges in realizing ethical decision-making in AVs. One of the salient challenges is that ethical dilemmas often involve complex trade-offs that are difficult to quantify and automate. For example, how should AVs decide in accident situations in which human harm is a likely consequence? and prioritize passenger safety over pedestrian safety? For example, the authors of [79] address ethical decision-making in complex urban driving scenarios in the case that human lives are at risk. They develop an ethical decision-making model based on DQN, and its main goal is to minimize injuries resulting from self-driving car crashes. The model considers factors such as the type of user involved in the accident, their age, and injury severity, the dynamic conditions of the self-driving car such as speed and acceleration, and the ethical considerations for the car. The authors of [80] address ethical decision-making during crashes involving humans and propose a predictive control framework for ethical decision-making in AD using rational ethics. The Lexicographic Optimization-based model predictive controller (LO-MPC) is designed to facilitate ethical rules implementation, In which obstacles and constraints are prioritized.
Different approaches can be taken for complex situations but still AVs are facing with the risk of accidents. The authors of [81] delegate the user/passenger the task of deciding what ethical approach should be taken by AVs in unavoidable accident scenarios. Indeed, passengers are enabled to ethically customize their AVs, namely, to choose between different settings corresponding to different ethical approaches or principles. The authors of [82] develop a computational model for building ethical AVs by learning and generalizing from human ethical judgments. They cast the problem of ethical learning for AVs as learning how to weigh the different features of the dilemma using utility calculus, to make these trade-offs reflect how people make them in a wide variety of ethical dilemmas. The authors of [83] investigate the ethics of AD and conclude that AD stakeholders are aware of the ethics of AD technology, and have different approaches concerning the authority of remote operators; the scientific community prioritizes safety and cybersecurity and agrees that AVs will not eliminate the risk of accidents.
@@ V-B Literature Review
Code of ethics need to be developed for AVs, and scientific bodies should provide recommendation. The German Ethics Commission [84] presents a code of ethics consisting of 20 ethical guidelines for Automated and Connected Driving and prioritizes safeguarding human life and rejecting discrimination based on personal traits in accidents. IEEE advocates for fair treatment and non-discrimination based on personal characteristics [85]. AVs often operate in dynamic environments where ethical considerations can change quickly. The system must be able to adapt its ethical decision-making processes in real-time. It must be ensured that the ML algorithms of AVs are transparent in their decision-making processes and that there is accountability for their actions. This involves not only technical challenges but also legal and regulatory considerations [86].To effectively address the above challenges and overcome the limitations of ML in the ethical assessment of AVs, it is important to bring humans into the loop to improve self-recognition and scene-recognition capabilities. Self-recognition refers to the AI’s ability to understand its state, limitations, and role in a given context. Scene recognition is about accurately interpreting the environment, including dynamic elements such as people, obstacles, and changes in the environment [89]. Table X summarizes the above findings.
V-C Key Findings and Insights
Ethical principles must be embedded in AVs to align their behavior with societal values and norms. To this end, ethical theories such as utilitarianism and deontology must be integrated into the decision-making processes of AVs. However, ethical decision-making in AVs faces several challenges. One major challenge is the complexity of ethical dilemmas that require difficult trade-offs. For example, in urban driving scenarios that aim to minimize injuries in accidents. Another challenge is the need to adapt ethical decision-making processes in real time. Ensuring transparency and accountability in these processes is also crucial. The ML algorithms used in AVs must be transparent in their decision-making and there must be accountability for their actions. One of the practical implementations of ethical decision-making in AVs is that users can customize the ethical settings of their vehicles. The inclusion of ethical guidelines and regulatory standards is essential. The German Ethics Commission and the IEEE, for example, advocate fair treatment and non-discrimination in the use of AVs. These guidelines help to standardize ethical practices in the development and use of AVs.
Stakeholders in AD are aware of the ethical issues surrounding AD technology, with varying opinions on remote operator authority. The scientific community prioritizes safety and cybersecurity but acknowledges that AVs will not eliminate accident risks. Formalizing safety and compatibility processes for AI and ML systems can enhance trust. However, ethical decision-making in AVs remains challenging due to the complexity of quantifying and automating ethical dilemmas, such as prioritizing passenger vs. pedestrian safety during accidents. An ethical decision-making model based on Deep Q-Network (DQN) is developed to minimize injuries in self-driving car crashes, considering factors like user type, age, injury severity, and vehicle dynamics.A computational model is developed to build ethical AVs by learning and generalizing from human ethical judgments. This involves using utility calculus to weigh different features of ethical dilemmas and reflect human decision-making in a variety of scenarios. Effective ethical assessment of AVs requires addressing both technical and regulatory challenges, including real-time adaptation of ethical decision-making and transparency in ML algorithms.
HITL-ML can improve self-recognition and scene recognition in AVs to assist in managing complex situations such as accidents and making ethical decisions in real-time. For example, in HITL-RL, human operators can help the RL algorithm recognize when its actions deviate from expected or safe behaviors and inject new actions. This helps the RL learn to better recognize and correct similar errors in the future. For scene recognition in AVs, the ML algorithm can ask human annotators to label complex or ambiguous data, such as those involving occlusions or unusual lighting conditions, ensuring that the ML learns from the most challenging examples.

### v2

@@ III-B2 Integration of Human Feedback in HITL-RL
III-B3 Agent-Specific vs Agent-Agnostic Protocols
Agent-specific protocols, meaning protocols for human interaction that are designed for a specific HITL-RL algorithm are preferred over agent-agnostic protocols. Agent-specific protocols allow humans to provide feedback and advice that is directly relevant to the AV learning algorithm. By understanding the vehicle specific learning process and representation, the protocols enable humans to ask more precise and informative questions. This leads to higher-quality guidance that is more likely to accelerate the vehicle’s learning and improve its decision-making capabilities. In contrast, agent-agnostic protocols are not specialized to the agent and therefore cannot ask human informative questions or have an observation model that accurately represents the human’s process for generating advice. Similarly, humans cannot provide optimally informative advice to the agent because they are unaware of the agent’s prior knowledge, exploration techniques, representation, or learning method[79].
III-B4 Evaluating HITL-ML Algorithms
Despite of significant contributions towards HITL-ML algorithms, most studies focus on improving the performance of algorithms rather than the interaction experience, but proper evaluation is essential to create practical applications for different types of end-users. The HITL-ML necessitates incorporating subjective user evaluations and acknowledging the cyclic influence between algorithms and users, making HITL-ML performance evaluation inherently subjective and complex [80][81].
III-B5 Robustness Against Adversarial Attacks
HITL-ML can make the models more robust against adversarial attacks. Adversarial attacks are deliberate attempts to fool ML models by providing deceptive input. When human are engaged in the model development process, they can provide valuable insights, detect anomalies, and make adjustments that might not be apparent through automated processes alone. This human involvement can help identify and mitigate potential vulnerabilities that adversarial attacks might exploit, thus reducing the severity and impact of such attacks on the models[82][83].
III-B6 Combining with safe RL Techniques
The addition of Safe RL techniques to HITL-RL setting can further reduce the likely of dangerous actions. Safe RL techniques ensure good performance and respect safety constraints during learning and deployment[82][84].
@@ IV Active Learning
AL optimizes the annotation process, accelerates development, and enhances system robustness by leveraging human expertise effectively. This can be very valuable in the scope of AVs, from UAV anomaly mitigation and semantic aerial mapping to intrusion detection and 3D object detection in AD. Annotating extensive data for training AVs is challenging due to their complex environments and unexpected situations. AL addresses these issues by reducing manual labeling through targeted instance selection, improving the quality of labeled data, and the accuracy of networks. The amount of data needed to train autonomous aerial or ground vehicles is enormous. AVs must navigate a variety of complex and diverse environments, which necessitate being represented in the training data. Moreover, AVs must deal with unexpected situations. Thus, to master these events, the model needs extensive data that captures such occurrences. Annotating such a volume of data for training autonomous aerial and ground vehicles can be an overwhelming and tedious task for humans and can pose a major challenge to AV development.
AL can be applied to this process, to cut down on the time and cost spent on training, while also increasing the accuracy of the networks. AL makes it possible to automate the selection process while choosing valuable data points. It starts by training a dedicated ML model on already-labeled data. The network then sorts through unlabeled data, uses the acquisition function to identify the instances it is most uncertain about, thereby finding data that would be challenging to the AVs algorithm and asks a human to label these specific instances which are added to training data and then the model is trained with labeled data. This approach reduces the number of instances that require manual labeling and ensures that the labeled data is as informative as possible [85].
Figure 4: AVs face unknown and unexpected scenarios and utilize AL to annotate them. The bottom right shows the AL loop comprising the steps of training, object selection, annotating, and appending. In the annotation step, human experts annotate the selected examples. AL impacts on unseen situations and training are shown in the top right.
@@
As can be seen in Fig. 4, AVs can be confronted with unknown and unexpected scenarios and AL can be used to annotate them. The bottom right of the figure shows the AL loop including the training, object selection, annotation and appending steps. In the annotation step, the selected examples are annotated by human experts. AL can help to cope with unseen situations, reducing training time and costs and improve training efficiency.
IV-A Literature Review
IV-A1 Security and Navigation
Pan etal. [46] aim to mitigate in-flight UAV anomalies by using a semi-supervised SVM. Since the pure semi-supervised SVM classification method may not classify well due to a lack of unlabeled data with rich content. AL is used to search the samples with rich information and significant differences for labeling to improve the quality of labeled data. Building on this concept, Ruckin etal. [47] have developed a new path planning framework for AL in aerial semantic mapping to improve UAV semantic awareness with minimal expert guidance. In an innovative approach, path planning goals are combined with AL features that allow us to adapt the UAV path to regions with informative training data. Similarly, Zhang etal. [48] design the intrusion detection system to detect unknown network attacks in the UAV network. The Open-CNN model is proposed to detect unknown attacks. Meanwhile, an AL approach is proposed for the unknown attacks based on the least confidence strategy, which enables the intrusion detection model to efficiently learn about observed novel network attacks with a low labeling budget. Table VIII summarizes the above findings.
@@ IV-A1 Security and Navigation
IV-A2 Detection
Liang etal. [52] are concerned with 3D object detection in AV. DL-based object detectors require large annotated datasets, which incurs compilation costs. To address this problem, they developed diversity-based AL, which uses the multimodal information of an AV dataset to propose a novel detection function that enforces spatial and temporal diversity in the selected samples. Simply, the most informative frames and objects are automatically selected for annotation by humans. Similarly, Ding etal. [21] address vehicle type recognition (VTR) from surveillance data in intelligent transportation. To facilitate manual labeling of large-scale surveillance data, they propose a VTR system based on deep AL. A deep neural network is efficiently trained by actively, incrementally, and automatically selecting unlabeled data points for human labeling. The training procedure is repeated several times so that the neural network is re-trained with the incrementally updated training data. Unlike traditional AL where the samples with high entropy are selected as the network output, in the proposed method, samples with both high and low entropy values are selected to be included in the training data. The proposed method is evaluated using the surveillance images from the CompCars public dataset. In addition, Satzoda etal. [86] develop a two-part technique for vehicle detection using an AL system and a symmetry-based iterative analysis. A modified AL technique is proposed that automatically selects positive and negative samples from multiple parts to train adaboost classifiers. The results of the classifiers are linked using a novel iterative window search algorithm and a symmetry-based regression model to extract fully visible vehicles. The AL method combined with the symmetry-based analysis has been shown to improve detection rates and also reduce false alarms compared to existing methods. Table IX summarizes the above findings.
IV-B Key Findings and Insights
From the survey in the previous subsection we identify the following open issues that need further research for better leveraging the AL potential in the scope of AVs:
IV-B1 Batch Training vs Sequential Training
@@ IV-B4 Uncertainty Sampling
IV-B5 Diverse Road Condition
@@ V Ethical Implications of AVs
We have seen that CL, HITL-RL and AL improve the capabilities of AVs. We emphasize that these HITL-ML techniques can promote self-recognition and scene recognition in AVs to assist in managing complex situations such as accidents, but ML algorithms in AI in general bring forth critical ethical and legal considerations. At the end, AVs should act as ethically rational agents and consider the ethical and legal implications of their actions, not just the outcomes.
Therefore, ethical principles must be encoded into AVs to ensure that their behavior aligns with societal values and norms. To this end, ethical theories and principles must be integrated into the decision-making processes of AVs. For example, an autonomous UAV can be considered a rational ethical agent if it accurately perceives its environment and makes decisions based on ethical frameworks such as utilitarianism where the ethical worth of an action is evaluated based on its consequences, with a focus on maximizing utility or happiness[87], or deontology where the ethical actions should be guided by a set of ethical principles or rules, regardless of the consequences [88], and acts in a way that is consistent with ethical principles, ensuring that its actions are “right” and beneficial to all stakeholders. The following subsection reviews the state-of-the-art in this topic.
V-A Literature Review
Complex situations involving human live need ethical consideration. Konert etal. [89] analyze the legal and ethical aspects of using UAVs, and state that we can achieve a higher degree of trust by formalizing the process of ensuring the safe and compatible functioning of systems based on AI and ML. However, there are some challenges in realizing ethical decision-making in AVs. One of the salient challenges is that ethical dilemmas often involve complex trade-offs that are difficult to quantify and automate. For example, how should AVs decide in accident situations in which human harm is a likely consequence? and prioritize passenger safety over pedestrian safety? For example, Vakili etal. [90] address ethical decision-making in complex urban driving scenarios in the case that human lives are at risk. They develop an ethical decision-making model based on DQN, and its main goal is to minimize injuries resulting from self-driving car crashes. The model considers factors such as the type of user involved in the accident, their age, and injury severity, the dynamic conditions of the self-driving car such as speed and acceleration, and the ethical considerations for the car. Wang etal. [91] address ethical decision-making during crashes involving humans and propose a predictive control framework for ethical decision-making in AD using rational ethics. The Lexicographic Optimization-based model predictive controller (LO-MPC) is designed to facilitate ethical rules implementation, In which obstacles and constraints are prioritized.
Different approaches can be taken for complex situations but still AVs are facing with the risk of accidents. Contissa etal. [92] delegate to the user/passenger the task of deciding which ethical approach should be taken by AVs in unavoidable accident scenarios. Indeed, passengers are enabled to ethically customize their AVs, namely, to choose between different settings corresponding to different ethical approaches or principles. Kim etal. [93] develop a computational model for building ethical AVs by learning and generalizing from human ethical judgments. They cast the problem of ethical learning for AVs as learning how to weigh the different features of the dilemma using utility calculus, to make these trade-offs reflect how people make them in a wide variety of ethical dilemmas. Martinho etal. [94] investigate the ethics of AD and conclude that AD stakeholders are aware of the ethics of AD technology, and have different approaches concerning the authority of remote operators; the scientific community prioritizes safety and cybersecurity and agrees that AVs will not eliminate the risk of accidents.
@@ V-A Literature Review
A code of ethics needs to be developed for AVs, and scientific bodies should provide recommendation. The German Ethics Commission [95] presents a code of ethics consisting of 20 ethical guidelines for Automated and Connected Driving and prioritizes safeguarding human life and rejecting discrimination based on personal traits in accidents. IEEE advocates for fair treatment and non-discrimination based on personal characteristics [96]. AVs often operate in dynamic environments where ethical considerations can change quickly. The system must be able to adapt its ethical decision-making processes in real-time. It must be ensured that the ML algorithms of AVs are transparent in their decision-making processes and that there is accountability for their actions. This involves not only technical challenges but also legal and regulatory considerations [97]. Table X summarizes the above findings.
V-B Key Findings and Insights
To effectively address the above challenges and overcome the limitations of ML in the ethical assessment of AVs, it is important to bring humans into the loop to improve self-recognition and scene-recognition capabilities. Self-recognition refers to the AI’s ability to understand its state, limitations, and role in a given context. Scene recognition is about accurately interpreting the environment, including dynamic elements such as people, obstacles, and changes in the environment [98].